# MSUHEP-24-019

### v1

@@ Abstract
The Hessian method is widely used in the global analysis of parton distribution functions (PDFs). It uses a set of orthogonal eigenvectors to calculate the first-order derivative terms of a physics observable, so that the leading order PDF-induced uncertainty can be estimated accordingly. In this article, we report an improved Hessian method which introduces the second-order derivatives in the uncertainty estimation. To estimate the effect of the second-order uncertainties, a pseudo global analysis is designed. We also find that the PDF-induced uncertainties on physics variables can be significantly different by taking the second-order effects into account, particularly when the precisions of experimental inputs used in the PDF global analysis are limited.
I Introduction
@@
In the past decades, a global analysis method based on quantum chromo-dynamics (QCD) has been established to study the proton structure. It sufficiently analyzes comprehensive data results from the deep inelastic scattering experiments with the energy scale of a few GeV, to the processes at hadron colliders at $\mathcal{O}(100)$ GeV. In practice, the global analysis is required to not only give the central predictions on parton densities corresponding to the best fit of the data, but also provide a way that the uncertainties can be easily calculated and extrapolated for any particular physics observable. The Hessian method [1] is one of the most widely used strategies to estimate the PDF uncertainties, such as in the latest CT18NNLO and MSHT20 PDFs [2, 3] which use both the DIS and hadron collider data. In this approach, a group of error PDF sets are provided, which represent the possible variations on the parameters of the non-perturbative functions of the quarks and gluons. These error sets correspond to orthogonal eigenvectors in the original parameter space, so that the uncertainty of a particular physics observable can be calculated using the error sets without complicated analysis on the correlations between the original parameters.
However, in the current Hessian method, only the leading order terms in the expansion with respect to a physics observable of those eigenvectors are taken into account to estimate the PDF-induced errors. As more data inputs are included in the global analysis, and more parameters are used in the non-perturbative parameterization of the PDFs, the higher order terms in the expansion of a physics observable may become more important. This has been mentioned in a few previous studies. For example, in the Error PDF Updating Method Package [4], contributions from the second-order terms in the expansion of a single eigenvector, appearing as the diagonal terms in the Hessian matrix as to be discussed later, were estimated based on the CT14 PDFs [5]. Based on the same PDFs, it was found later in Ref. [6] that the higher order terms of a single eigenvector can be sizable in the analysis of the CMS di-jet measurement. However, the off-diagonal terms corresponding to the expansion of multiple eigenvectors cannot be faithfully tested in the above studies since CT14 did not provide the needed error sets for this kind of tasks.
@@
In this work, we introduce an improved Hessian method, which generated the PDF error sets allowing a complete second-order calculation of the uncertainty of a physics observable. To make a detailed comparison between the improved method and the original one, a pseudo global fit is studied. As will be demonstrated in this test, when including the second-order terms, the uncertainty on parton densities could be noticeably different from the original leading order estimation. Such difference can be reduced as the precision of the input data in the global analysis improves. However, the required precision is beyond many experimental inputs included in the present PDF global analysis, especially from those old DIS and fixed-target Drell-Yan data [2]. Hence, it would be important to consider the higher order terms in the uncertainty estimation of the PDF global analysis.
II Method
The Hessian method [1] is designed based on the following expansion:
$\mathcal{O}(z_{i})=\mathcal{O}(z^{0}_{i})+\sum_{i}\frac{\partial\mathcal{O}}{\partial z_{i}}\Big{|}_{z^{0}_{i}}\text{d}z_{i}+\frac{1}{2}\sum_{i,j}\frac{\partial^{2}\mathcal{O}}{\partial z_{i}\partial z_{j}}\Big{|}_{z^{0}_{i},z^{0}_{j}}\text{d}z_{i}\text{d}z_{j}+\cdot\cdot\cdot$ (1)
where $\mathcal{O}$ is a physical observable whose theory prediction depends on PDFs. $\{z_{i}\}$ represent the parameters in the PDFs, and $\{z^{0}_{i}\}$ are the corresponding central values determined as the best fit to the data inputs. When $z_{i}$ is independent of $z_{j}$, for $i\neq j$, the leading order uncertainty on $\mathcal{O}$ can simply be calculated as:
$\delta\mathcal{O}\approx\sqrt{\sum_{i}\left(\frac{\mathcal{O}(z^{0}_{i}+\delta z_{i})-\mathcal{O}(z^{0}_{i}-\delta z_{i})}{2}\right)^{2}}$ (2)
where $\delta z_{i}$ is the uncertainty on $z_{i}$. $\mathcal{O}(z^{0}_{i}\pm\delta z_{i})$ are the calculations of $\mathcal{O}$ where the value of the $i$-th parameter is shifted by ±$\pm\delta z_{i}$, while all other parameters $z_{j}(j\neq i)$ are fixed at their central values $z^{0}_{j}$. To enable the calculation, the global analysis provides a central PDF set $S_{0}=S(z^{0}_{i})$ which can be used to calculate the central prediction of $\mathcal{O}(S_{0})$, and a group of error PDF sets $S_{i\pm}=S(z^{0}_{i}\pm\delta z_{i};z_{j}=z^{0}_{j})$ for $j\neq i$. In this strategy, the first-order derivatives are represented by
$\frac{\partial\mathcal{O}}{\partial z_{i}}\Big{|}_{z^{0}_{i}}=\frac{\mathcal{O}(z^{0}_{i}+\delta z_{i})-\mathcal{O}(z^{0}_{i}-\delta z_{i})}{2\delta z_{i}},$ (3)
while the higher order derivatives $\partial^{2}\mathcal{O}/\partial z_{i}\partial z_{j}$ are ignored. As discussed above, Ref. [4] and Ref. [6] studied effects on $\partial^{2}\mathcal{O}/\partial z^{2}_{i}$.
@@ II Method
We first review how to obtain the orthogonal eigenvectors $\{v_{i}\}$, with $\{z_{i}\}$ being the displacement (in the space of $\{z_{i}\}$) along the direction of $\{v_{i}\}$, by the Hessian method. The non-perturbative functions which describe the parton densities at the scale of $Q_{0}$, taken from the CT18NNLO PDF as an example, are written in the formalism as:
@@
$\chi^{2}=\sum_{\alpha}\frac{[D_{\alpha}-T_{\alpha}(a_{i})]^{2}}{\sigma^{2}_{\alpha}}$ (5)
where $D_{\alpha}$ and $\sigma_{\alpha}$ are the experimental inputs and their corresponding uncertainties. $T_{\alpha}(a_{i})$ is the theory prediction of $D_{\alpha}$, depending on the PDF shape parameters $\{a_{i}\}$. In practice, $\chi^{2}_{\text{total}}$ is more complicated, which may include correlated systematic uncertainties. Furthermore, to deal with the potential biases of the theory calculation and the experimental measurements, possible inconsistency between data, and the uncertainty due to the choice of the non-perturbative parameterizations, some dynamical and/or global tolerances are also introduced to define the errors of the fitted PDF parameters, such as in the CT18 and MSHT20 [2, 3]. For simplicity, we ignore all these variants in this work as they will not strongly alter our conclusions. In this work, we treat the pseudo-PDF uncertainty to be uncorrelated and purely statistical.
Given the fact that the best fit values of $\{a_{i}\}$ appear near the minimal $\chi^{2}$, we have:
@@
The Hessian matrix is then defined as $H_{ij}=\partial^{2}\chi^{2}/\partial a_{i}\partial a_{j}$. Note that the original parameters $a_{i}$ can be correlated in the global fit because experimental observables usually relate to multiple PDF shape parameters. One can convert $\{a_{i}\}$ to another set of orthogonal parameters $\{z_{i}\}$ by diagonalizing $H_{ij}$:
@@
where the scaling factor $s_{i}$ is introduced to normalize $z_{i}$. Usually, $s_{k}$ is chosen to make
@@
so that the best fit of $\chi^{2}$ (denoted as $\chi^{2}_{\text{min}}$) corresponds to $\{z_{i}=0\}$, and the one-standard deviation uncertainty in the direction of the $i$-th eigenvector corresponds to $\{z_{i}=\pm 1;z_{j}=0$ for $j\neq i\}$. By this requirement, $s_{i}$ are approximately equal to $\sqrt{2/\epsilon_{i}}$.
The central PDF set and the error sets can be established. The central PDF set corresponds to the best fit parameters $S_{0}=S(a_{i}=a^{0}_{i})$, or $S_{0}=S(z_{i}=0)$. For the error PDF sets, we denote the ±$\pm 1$ standard variation in the direction of the $i$-th eigenvector as
$S_{i\pm}=S(z_{i}=\pm 1;z_{j}=0(j\neq i)).$ (9)
Hence, the derivative in Eq. (3) of any physical observable $\mathcal{O}$ can be calculated using the PDF error sets:
$\frac{\partial\mathcal{O}}{\partial z_{i}}\Big{|}_{0}=\frac{\mathcal{O}(S_{i+})-\mathcal{O}(S_{i-})}{2},$ (10)
so that the leading order uncertainty is acquired. $S_{0}$ and $S_{i\pm}$ are the PDF sets provided by the global analysis groups such as CTEQ-TEA and MSHT. They can be used to calculate the first-order derivatives in Eq. (1), but not those off-diagonal terms.
Below, we show how to obtain the PDF error sets for calculating the full second-order derivatives in Eq. (1). Following the definition of the derivative, the needed error sets are:
@@
in which two displacements are varied simultaneously, so that the second-order derivatives in Eq. (3) can be calculated as:
@@
Note that for $i=j$, the second-order error sets are $S_{i++}=S(z_{i}=2;z_{k}=0$ for $k\neq i)$ and $S_{i--}=S(z_{i}=-2;z_{k}=0$ for $k\neq i)$. Consequently, we have
@@
According to Eq. (II), the values of the original parameters $\{a_{i}\}$ corresponding to $\{z_{i}=\pm 1,z_{j}=\pm 1;z_{k}=0$ for $k\neq i,j\}$ can be calculated, and thus the second-order error sets $S^{i\pm}_{j\pm}$ can be directly generated. Since the Hessian matrix comes from the derivatives of $\chi^{2}$, the second-order error sets can only be acquired together with the first-order error sets in a complete PDF global analysis.
With full second-order derivatives included, the uncertainty on $\mathcal{O}$ is written as:
@@
where the first, second and third terms correspond to the contributions from the first-order derivatives, the second-order diagonal derivatives, and the second-order off-diagonal derivatives in Eq. (1). The process of proof of Eq. (14) is given in the Appendix. One can also calculate the uncertainty in an alternative way:
@@
The bootstrap procedure gives consistent uncertainty with Eq. (14). It is also convenient to calculate the correlation of the PDF-induced uncertainties between two physical observables $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$. One can calculate $(\delta\mathcal{O}_{1})^{2}$, $(\delta\mathcal{O}_{2})^{2}$ and $(\delta(\mathcal{O}_{1}+\mathcal{O}_{2}))^{2}$ via either aforementioned method. The correlation is then easily acquired on the basis of its original definition of correlation:
@@
III Pseudo global analysis
@@
A group of pseudo observables are also constructed according to the PDFs in Eq. (III) and Eq. (III). These observables, with uncertainties assigned, will be used as the data input in the pseudo global analysis. First, we introduce $\sigma^{\text{DIS}}_{\gamma}(x)$ and $\sigma^{\text{DIS}}_{Z}(x)$:
@@
as pseudo observables similar to the cross section measurements of the photon-exchange DIS process $\ell+p\rightarrow\gamma^{*}\rightarrow\ell+X$, and the $Z$-exchange DIS process $\ell+p\rightarrow Z/\gamma^{*}\rightarrow\ell+X$. The $q(x)$ functions are taken from Eq. (III) and Eq. (III), with $\{a_{i}\}$ values from Tab. 1. The coefficient 0.25 (1.2) roughly reflects the strength difference between the down-type and up-type quarks coupling to $\gamma$ ($Z$). Note that we take the definitions of $u\equiv u_{V}+\bar{u}$ and $d\equiv d_{V}+\bar{d}$, so that the $\bar{u}$ and $\bar{d}$ terms have a coefficient of 2. For simplicity, we ignore other contributions in a real calculation of the DIS process, such as the evolution of PDFs as a function of $Q$, the electroweak couplings and the higher order corrections. Although these effects would significantly change the predictions on any physical observable, the main sensitivity to the particular parton information in those observables remains. For $\sigma^{\text{DIS}}_{\gamma}(x)$, 28 data points are calculated, with $x=$ 0.0001, 0.0002, 0.0003, 0.0004, 0.0005, 0.0006, 0.0007, 0.0008, 0.0009, 0.001, 0.002, 0.003, 0.004, 0.005, 0.006, 0.007, 0.008, 0.009, 0.01, 0.02, 0.03, 0.04, 0.05, 0.06, 0.07, 0.08, 0.09, 0.1. For $\sigma^{\text{DIS}}_{Z}(x)$, 16 points are calculated, with $x=$ 0.01, 0.02, 0.03, 0.04, 0.05, 0.06, 0.07, 0.08, 0.09, 0.1, 0.15, 0.2, 0.25, 0.3, 0.35, 0.4.
Similarly, we introduce the following cross section observables in the proton-proton collisions: $\sigma^{\text{pp}}_{Z}(x_{1},x_{2})$, $\sigma^{\text{pp}}_{W^{+}}(x_{1},x_{2})$ and $\sigma^{\text{pp}}_{W^{-}}(x_{1},x_{2})$ as:

### v2

@@ Abstract
The Hessian method is widely applied in the global analysis of parton distribution functions (PDFs), which uses a set of orthogonal eigenvectors to give predictions of a physical observable. Its uncertainty is estimated based on the assumption that all physical observables can be approximately expressed as linear functions of the non-perturbative parameters in PDF. In this article, we report an improved Hessian method which takes the non-linear effects into account in the uncertainty estimation. A pseudo global analysis is designed to numerically test the new method. The non-linear uncertainties can significantly enlarge the original linear ones. Such uncertainties can be reduced with high precision data introduced into the global analysis. However, the non-linear effect could still be sizable corresponding to the current precision of the experimental measurements.
I I. Introduction
@@
In the past decades, a global analysis method based on quantum chromodynamics (QCD) has been established to study the proton structure at high energy scale. It sufficiently analyzes comprehensive data results from the deep inelastic scattering (DIS) experiments with the energy scale of a few GeV, to the processes at hadron colliders at $\mathcal{O}(100)$ GeV. The global analysis is expected to not only give the central predictions on parton densities corresponding to the best fit of the data, but also provide a way that the uncertainties can be easily calculated and extrapolated for any particular physical observable. The Hessian method [1] is one of the most widely used strategies to estimate the PDF uncertainties, such as in the latest CT18NNLO and MSHT20 PDFs [2, 3]. In this approach, a group of error PDF sets are provided, which represent the possible variations on the parameters of the non-perturbative functions of the quarks and gluons. These error sets correspond to orthogonal eigenvectors in the original parameter space, so that the uncertainty of a particular physical observable can be calculated using the error sets without complicated analysis on the correlations between the original parameters.
In the current Hessian method, the uncertainty estimation is established as a linear approximation. Firstly, the orthogonal eigenvectors and their corresponding uncertainties are determined by assuming the physical observables of those data used in the global analysis can be expressed as linear functions of the non-perturbative PDF parameters. Secondly, when computing the PDF-induced uncertainty on a given physical observable using the eigenvectors, the observable is also assumed as linear function of the eigenvectors. The non-linear effects, appearing as the higher order derivatives in the expansion of physical observables as functions of PDF parameters or eigenvectors, are ignored. Such effect could be significant in high energy interactions, such as the LHC which provides precise experimental results in today’s PDF global analysis. A few previous studies have been performed on the non-linearities in the global analysis. For example, Ref. [4] discussed contributions from the second-order terms based on the CT14 PDFs [5], but only for those appearing as the diagonal elements in the Hessian matrix (as to be introduced later). Based on the same PDFs, it was found later in Ref. [6] that the diagonal higher order terms could be sizable in the analysis of the CMS di-jet measurement.
@@
In this work, we introduce an improved Hessian method, in which the non-linear effects are estimated as additional uncertainties. To make a detailed comparison between the improved method and the original one, a pseudo global fit is studied. As will be demonstrated in this test, the uncertainty on parton densities could be noticeably different from the original linear estimation. Such difference can be reduced as the precision of the input data in the global analysis improves. However, the needed precision to reduce the non-linear uncertainties to negligible level is far beyond the current precision of experimental observations.
The article will be arranged as following: In section II, we review the general strategy of Hessian method; In section III, we discuss the non-linear effects in the global analysis; In section IV, we introduce the improved method to estimate the non-linear uncertainties; In section V, we perform a pseudo global analysis study to numerically test the improved method, and summarize the conclusion of this article.
II II. The original Hessian method
@@
In this section, we review the original Hessian method [1] and its linear uncertainty estimation. In Hessian method, a series of non-perturbative functions are introduced to describe the parton densities at the scale of $Q_{0}$. Taking from the CT18NNLO PDF as an example, the functions are written in the formalism as:
@@
$\chi^{2}=\sum_{\alpha}\frac{[D_{\alpha}-T_{\alpha}(\{a_{i}\})]^{2}}{\sigma^{2}_{\alpha}}$ (2)
where $D_{\alpha}$ and $\sigma_{\alpha}$ are the experimental inputs and their corresponding uncertainties. $T_{\alpha}(a_{i})$ is the theory prediction of $D_{\alpha}$, depending on the PDF shape parameters $\{a_{i}\}$. Given the fact that the best fit values of $\{a_{i}\}$ appear near the minimal $\chi^{2}$, we have:
@@
The Hessian matrix is then defined as $H_{ij}=\partial^{2}\chi^{2}/\partial a_{i}\partial a_{j}$. Note that $\{a_{i}\}$ are quite independent as free parameters in the non-perturbative functions, but their uncertainties in a global analysis can be correlated because experimental observables usually relate to multiple PDF shape parameters. The correlation depends on the particular set of experimental inputs used in a global fit. One can convert $\{a_{i}\}$ to another set of orthogonal parameters $\{z_{i}\}$ by diagonalizing $H_{ij}$:
@@
where the scaling factors $\{s_{i}\}$ are introduced to normalize $\{z_{i}\}$. Usually, $\{s_{i}\}$ are chosen to make
@@
so that the best fit of $\chi^{2}$ (denoted as $\chi^{2}_{\text{min}}$) corresponds to $\{z_{i}=0\}$, and the one-standard deviation uncertainty in the direction of the $i$-th eigenvector corresponds to $\{z_{i}=\pm 1;z_{j}=0$ for $j\neq i\}$. By this requirement, $\{s_{i}\}$ are approximately equal to $\sqrt{2/\epsilon_{i}}$.
In practice, $\chi^{2}$ is more complicated, which may include correlated systematical uncertainties. Furthermore, the one-standard deviation uncertainty is usually enlarged by some dynamical and/or global tolerances to deal with the possible inconsistency between data, and to cover the uncertainty due to the choice of the non-perturbative parameterizations. For simplicity, we ignore all these variants in this work as they will not strongly alter our conclusions.
For a given physical observable whose theory prediction depends on PDFs, it can be expanded as:
$\mathcal{O}(\{z_{i}\})=\mathcal{O}(\{z^{0}_{i}\})+\sum_{i}\frac{\partial\mathcal{O}}{\partial z_{i}}\Big{|}_{z^{0}_{i}}\text{d}z_{i}+\frac{1}{2}\sum_{i,j}\frac{\partial^{2}\mathcal{O}}{\partial z_{i}\partial z_{j}}\Big{|}_{z^{0}_{i},z^{0}_{j}}\text{d}z_{i}\text{d}z_{j}+\cdot\cdot\cdot$ (6)
As $\{z_{i}\}$ represent orthogonal eigenvectors, the leading order uncertainty on $\mathcal{O}$ can simply be calculated as:
$\delta\mathcal{O}$ $\approx$ $\sqrt{\sum_{i}\left[\frac{\mathcal{O}(z^{0}_{i}+\delta z_{i})-\mathcal{O}(z^{0}_{i}-\delta z_{i})}{2\delta z_{i}}\right]^{2}}$ (7)
where $\delta z_{i}=1$ is the uncertainty on $z_{i}$. $\mathcal{O}(z^{0}_{i}\pm\delta z_{i})$ is the calculation of $\mathcal{O}$ where the value of the $i$-th parameter is shifted by $\pm\delta z_{i}$, while all other parameters $z_{j}(j\neq i)$ are fixed at their central values $z^{0}_{j}$. To enable the calculation, the global analysis provides a central PDF set $S_{0}=S(\{z^{0}_{i}\})$ which can be used to calculate the central prediction of $\mathcal{O}(S_{0})$, and a group of error PDF sets $S_{i\pm}=S(z^{0}_{i}\pm\delta z_{i};z_{j}=z^{0}_{j})$ for $j\neq i$. Hence, the uncertainty in Eq. (7) can be calculated using the PDF error sets:
$\delta\mathcal{O}=\sqrt{\sum_{i}\left(\frac{\mathcal{O}(S_{i+})-\mathcal{O}(S_{i-})}{2}\right)^{2}}$ (8)
III III. Non-linear effects in global analysis
The first linear assumption in the original Hessian method is introduced in the determination of the eigenvectors and their uncertainties. The orthogonal parameters $\{z_{i}\}$ are acquired using the Hessian matrix containing only the $\partial^{2}\chi^{2}/\partial a_{i}\partial a_{j}$ terms. The one standard deviation is defined as $\delta z_{i}=1$ corresponding to $\Delta\chi^{2}=1$. All these settings are based on the assumption that the physical observables of the data used in the global analysis can be written as the linear functions of the original PDF parameters $\{a_{i}\}$:
$T_{\alpha}(\{a_{i}\})=\sum_{i}t^{i}_{\alpha}a_{i}$ (9)
where $t^{i}_{\alpha}$ is the coefficient. In principle, Eq. (9) never holds true in global analysis. Firstly, the non-perturbative functions in Eq. (1) are naturally non-linear, e.g. the power terms $x^{1-a_{1}}$ and $(1-x)^{a_{2}}$ which are guided by the Regge theory [7] and the spectator counting rules [8] to describe the asymptotic behavior of $f_{q}(x)$ in the limits $x\rightarrow 0$ and 1. Secondly, the calculations on physical interactions are usually convolutions over $x$. Thirdly, for those collider and fixed target experiments, the calculations on the two-hadron initial state contain the product of the two parton densities. In practice, when data inputs used in a global analysis have high precisions, $\Delta\chi^{2}$ would increase rapidly as $\{a_{i}\}$ values running away from their best fitted values. Thus, all calculations can be done in small intervals around $\{a^{0}_{i}\}$ so that $T_{\alpha}(\{a_{i}\})$ can be numerically linear. In other words, the non-linearity depends on the precision of the experimental measurements in the PDF global analysis. In current PDFs, the experimental inputs provide good constraints in some $x$ regions and for the quark flavors that they are sensitive to. However, the data precisions in the foreseeable future would not be good enough to suppress the non-linear effect for all quark flavors and in a wide range of $x$.
Another linear assumption is introduced in computing the PDF-induced uncertainties on physical observables using the eigenvectors. In the current Hessian method, the higher order contributions in Eq. (6) are ignored when extrapolating the uncertainties on eigenvectors to a given physical observable. The Hessian-type PDFs such as CT18 and MSHT20 provide only the error sets $S_{i\pm}$ corresponding to the complete differential of $z_{i}$, which allows calculation of $\partial\mathcal{O}/\partial z_{i}$, $\partial^{2}\mathcal{O}/\partial z^{2}_{i}$ or even higher order contributions as the diagonal elements. But these error sets cannot give estimations on the non-linear terms such as $\partial^{2}\mathcal{O}/\partial z_{i}\partial z_{j}$. The higher order contributions can be significant especially when the computed physical observable corresponds to $x$ region or quark flavor very differently from that of the data in the global fit.
In CT18 [2], the uncertainties of the PDF error sets are quoted to cover the difference from varying the formalism of the non-perturbative functions of Eq. (1). By doing this, CT18 gives more conservative uncertainty estimation compared to MSHT20 [3] and NNPDF4 [10], which use similar data set in the global fit. The enlarged uncertainties reflect the non-linear effects in global analysis, because if the non-perturbative functions tested by CT18 were all in good linear approximation, they would lead to consistent predictions on any physical observable because they were connected by linear transformations in the PDF parameter space. In other words, it is, to some extent, similar to estimating the uncertainty due to the non-linearity with respect to the uncertainty due to the particular choice of non-perturbative functions used in the global fit. However, it is inevitable in the enumeration method that many functions being tested are known to have disadvantages in describing the data. Consequently, the uncertainty may be overestimated by quoting the imperfection of the excluded functions rather than the imperfection of the chosen function. On the other hand, the uncertainty may also be underestimated as there is always possible functions missed in the enumeration. Therefore, a new method to estimate the uncertainty of the chosen function itself would be important.
IV IV. Improved Hessian method
The improved method will be derived in this Section. Then, in Section V we will apply it to a pseudo analysis, to test how non-linear contributions can affect a global analysis. The pseudo-linear calculation will be compared to the nominal (un-improved) global fit.
The strategy is as following:
i) For a given non-perturbative function formalism such as in Eq. (1), perform the global analysis with the normal method, giving the best fitted values of $\{a_{i}=a^{0}_{i}\}$ and the central PDF set $S_{0}$. In this nominal fit, $T_{\alpha}(\{a_{i}\})$ are predicted with different $\{a_{i}\}$ inputs based on the conventional calculations.
ii) For each physical observable, a set of linear factors are defined as:
$K^{\alpha}_{i}$ $=$ $\frac{T^{i+}_{\alpha}-T^{i-}_{\alpha}}{2\epsilon_{i}}.$ (10)
where $T^{i+}_{\alpha}=T_{\alpha}(\{a_{i}=a^{0}_{i}+\epsilon_{i};a_{j}=a^{0}_{j}\text{ for }i\neq j\})$ and $T^{i-}_{\alpha}=T_{\alpha}(\{a_{i}=a^{0}_{i}-\epsilon_{i};a_{j}=a^{0}_{j}\text{ for }i\neq j\})$. Apparently, $K^{\alpha}_{i}$ is an approximation of $\partial T_{\alpha}(\{a_{i}\})/\partial a_{i}$ around $\{a_{i}=a^{0}_{i}\}$ in the range of $(a_{i}-\epsilon_{i},a_{i}+\epsilon_{i})$. It represents the slope of the secant line, as shown in Fig. 1. The $\epsilon_{i}$ parameter corresponds to the range of $a_{i}$ in which the linear derivation of Eq. (10) is assumed. Thus, $\epsilon_{i}$ can simply be set to the original uncertainty of the $a_{i}$ parameter given by the fit obtained form i).
Figure 1: The secant line of $T_{\alpha}(a_{i};a_{j}=a^{0}_{j}\text{ for }j\neq i)$ around the best fitted value $a^{0}_{i}$.
iii) A group of new observables can then be established as linear functions of $\{a_{i}\}$:
$T^{(r)}_{\alpha}(\{a_{i}\})=T_{\alpha}(\{a^{r}_{i}\})+\sum_{i}K^{\alpha}_{i}\times(a_{i}-a^{r}_{i}),$ (11)
where $\{a^{r}_{i}\}$, $r=0,1,\cdot\cdot\cdot,n$ represent a set of points in the parameter space of $\{a_{i}\}$, with $n$ as the number of the parameters. Values of $\{a^{r}_{i}\}$ also represent the range of the linear assumption, but in terms of $T_{\alpha}$. In order to achieve this, $\{a^{r}_{i}\}$ values can be set corresponding to $\{z_{r}=\Delta_{r};z_{i}=0\text{ for }i\neq r\}$. $\Delta_{r}$ is chosen to have $T_{\alpha}(\{a^{r}_{i}\})-T_{\alpha}(\{a^{0}_{i}\})$ consistent with $\sigma_{\alpha}$ for those data inputs sensitive to $z_{r}$. Note that when the eigenvector $z_{r}$ is constrained by multiple data inputs in the global fit, the uncertainty of $z_{r}$ ($\delta z_{r}$) will be smaller than $\Delta_{r}$, since $\delta z_{r}$ reflects the combination of the statistical power of those data inputs.
Using Eq. (11), one can perform a linear fit, in which the best values of $\{a_{i}\}$ are determined by requiring the minimum of $\chi^{2}(T^{(r)})$ defined as:
$\chi^{2}(T^{(r)})=\sum_{\alpha}\frac{[D_{\alpha}-T^{(r)}_{\alpha}(\{a_{i}\})]^{2}}{\sigma^{2}_{\alpha}}.$ (12)
When Eq. (9) is in good approximation, the linear fit using Eq. (12) would give best fit values, denoted as $\{a^{r,0}_{i}\}$ for each $r$, consistent with the nominal fit $\{a^{0}_{i}\}$. Therefore, the difference between the two fits reflects the non-linearity of the physical observables $T_{\alpha}(\{a_{i}\})$, and can be quoted as additional uncertainties. To do this, an additional PDF error set is defined as $S_{E}(r)=S(\{a^{r,0}_{i}\})$ for each eigenvector. For a given physical observable $\mathcal{O}$, the non-linear uncertainty is defined as:
$\delta\mathcal{O}_{E}=\sqrt{\sum_{r}\left[\mathcal{O}(S_{E}(r))-\mathcal{O}(S_{0})\right]^{2}}$ (13)
It should be emphasized that Eq. (13) quantifies the impact on non-linearity when extrapolated to a physical observable. Such non-linearity introduces a bias in the analysis, which may in turn contribute to the PDF uncertainty. The notions of “bias” and “uncertainty” are not strictly equivalent, although they are likely to be of comparable size. A more detailed analysis would be required to establish their precise relationship in a specific case. $\delta\mathcal{O}_{E}$ is the uncertainty of assuming $T_{\alpha}(\{a_{i}\})$ as linear functions, which is independent of how to extrapolate the original experimental uncertainties $\sigma_{\alpha}$ to each eigenvector, and further to a physical observable $\mathcal{O}$. As discussed in the section II, the error sets $S_{i\pm}$ allow calculations only on the first-order derivatives in Eq. (6). Below, we show how to obtain the PDF error sets for calculating the full second-order derivatives. Following the definition of the derivative, the needed error sets are:
@@
in which two displacements are varied simultaneously, so that the second-order derivatives can be calculated as:
@@
For $i=j$, the second-order error sets can be calculated as $S_{i++}=S(z_{i}=2;z_{k}=0$ for $k\neq i)$ and $S_{i--}=S(z_{i}=-2;z_{k}=0$ for $k\neq i)$. Consequently, we have
@@
According to Eq. (II), the values of the original parameters $\{a_{i}\}$ corresponding to $\{z_{i}=\pm 1,z_{j}=\pm 1;z_{k}=0$ for $k\neq i,j\}$ can be calculated, and thus the second-order error sets $S^{i\pm}_{j\pm}$ can be directly generated. Since the Hessian matrix comes from the derivatives of $\chi^{2}$, the second-order error sets can only be acquired together with the first-order error sets in a complete PDF global analysis. With the full second-order derivatives included, the uncertainty on $\mathcal{O}$ is written as:
@@
where the first, second and third terms correspond to the contributions from the first-order derivatives, the second-order diagonal derivatives, and the second-order off-diagonal derivatives in Eq. (6), respectively. The process of proof of Eq. (17) is given in the Appendix.
One can also calculate the uncertainty in an alternative way:
@@
The bootstrap procedure gives consistent uncertainty with Eq. (17). It is also convenient for calculating the correlation of the PDF-induced uncertainties between two physical observables $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$. One can calculate $(\delta\mathcal{O}_{1})^{2}$, $(\delta\mathcal{O}_{2})^{2}$ and $(\delta(\mathcal{O}_{1}+\mathcal{O}_{2}))^{2}$ via either aforementioned method. The correlation is then easily acquired on the basis of its original definition of correlation:
@@
Note that both the bootstrap method and Eq. (17) assume $\{z_{i}\}$ have independent gaussian distributions. As will be discussed in the next section, it is equivalent to the linear assumptions of Eq. (9), of which the uncertainty has been estimated as $\delta\mathcal{O}_{E}$. In conclusion, the total uncertainty on a physical observable $\mathcal{O}$ is the quadratic combination of the original first-order uncertainties, the second-order uncertainties, and the non-linear uncertainty of $\delta\mathcal{O}_{E}$:
$\left(\delta\mathcal{O}_{\text{total}}\right)^{2}=\frac{1}{4}\sum_{i}\left[\mathcal{O}(S_{i+})-\mathcal{O}(S_{i-})\right]^{2}$ (19)
$+$ $\frac{1}{32}\sum_{i}\left[\mathcal{O}(S_{i++})+\mathcal{O}(S_{i--})-2\mathcal{O}(S_{0})\right]^{2}$
$+$ $\frac{1}{32}\sum_{i\neq j}\left[\mathcal{O}(S^{i+}_{j+})-\mathcal{O}(S^{i-}_{j+})-\mathcal{O}(S^{i+}_{j-})+\mathcal{O}(S^{i-}_{j-})\right]^{2}$
$+$ $\sum_{r}\left[\mathcal{O}(S_{E}(r))-\mathcal{O}(S_{0})\right]^{2}.$
[…]
V V. Pseudo global analysis
@@
A group of pseudo observables are also constructed according to the PDFs in Eq. (V) and Eq. (V). These observables, with uncertainties assigned, will be used as the data input in the pseudo global analysis. First, we introduce $\sigma^{\text{DIS}}_{\gamma}(x)$ and $\sigma^{\text{DIS}}_{Z}(x)$:
@@
as pseudo observables similar to the cross section measurements of the photon-exchange DIS process $\ell+p\rightarrow\gamma^{*}\rightarrow\ell+X$, and the $Z$-exchange DIS process $\ell+p\rightarrow Z/\gamma^{*}\rightarrow\ell+X$. The $q(x)$ functions are taken from Eq. (V) and Eq. (V), with $\{a_{i}\}$ values from Tab. 1. The coefficient 0.25 (1.2) roughly reflects the strength difference between the down-type and up-type quarks coupling to $\gamma$ ($Z$). Note that we take the definitions of $u\equiv u_{V}+\bar{u}$ and $d\equiv d_{V}+\bar{d}$, so that the $\bar{u}$ and $\bar{d}$ terms have a coefficient of 2. For simplicity, we ignore other contributions in a real calculation of the DIS process, such as the evolution of PDFs as a function of $Q$, the electroweak couplings and the higher order corrections. Although these effects would significantly change the predictions on any physical observable, the main sensitivity to the particular parton information in those observables remains. For $\sigma^{\text{DIS}}_{\gamma}(x)$, 28 data points are calculated, with $x=$ 0.0001, 0.0002, 0.0003, 0.0004, 0.0005, 0.0006, 0.0007, 0.0008, 0.0009, 0.001, 0.002, 0.003, 0.004, 0.005, 0.006, 0.007, 0.008, 0.009, 0.01, 0.02, 0.03, 0.04, 0.05, 0.06, 0.07, 0.08, 0.09, 0.1. For $\sigma^{\text{DIS}}_{Z}(x)$, 16 points are calculated, with $x=$ 0.01, 0.02, 0.03, 0.04, 0.05, 0.06, 0.07, 0.08, 0.09, 0.1, 0.15, 0.2, 0.25, 0.3, 0.35, 0.4.
Similarly, we introduce the following cross section observables in proton-proton collisions: $\sigma^{\text{pp}}_{Z}(x_{1},x_{2})$, $\sigma^{\text{pp}}_{W^{+}}(x_{1},x_{2})$ and $\sigma^{\text{pp}}_{W^{-}}(x_{1},x_{2})$ as: